# Gemstones: A Model Suite for Multi-Faceted Scaling Laws

### v2

@@ Appendix H The Price of Stepping Off the Scaling Law
If You Value Your Time, Train Wide Models. We first show that in our training setup, training wider models is far more efficient than training deep models. In Figure˜16, we reflect on the consequences of suboptimal architectural choices, by considering how much of a given resource—FLOPs or GPU hours—would be “overspent” to reach any target loss value with the plotted architecture rather than the prescribed width and depth. We find that choosing to train “skinny” models (top left) wastes many FLOPs and GPU hours. The scale of overspend is quite different however, with the least efficient models only overspending about $50\%$ on FLOPs but wasting more than $200\%$ of the GPU hours spent by the best configuration. In other words, in the time taken to train a single (very) suboptimal model to the desired loss value, one could train three optimal-width-depth models. We note that while the time-optimal models tend to be the wider ones, this is probably due to our training scheme. Similar to other open-source efforts such as OLMo et al. [2024], we do not make any use of pipeline parallelism. In summary, for standard parallelism implementations, wider models are simply easier to scale, but as a result our observations regarding resource overspending may not generalize to other parallelism strategies.
Appendix I Scaling Laws Predict That Overtraining Is Efficient.
@@
Similarly to Gadre et al. [2024], we can shift optimal points to simulate overtraining. To do this, we fix a FLOP budget and trace out a path of model sizes and corresponding token counts to remain within that budget. For each model size and token count, we record the “overtraining factor,” which is the selected number of training tokens divided by the optimal number of tokens for that model shape. An overtraining factor of less than one corresponds to undertraining the model, and a factor greater than one represents overtraining. We show the results of this process in Figure˜17. We see that overtraining does increase predicted loss at a given FLOP count but that these curves are actually quite flat. We include the loss values of open source models on our own validation set to allow readers to contextualize the y-axis values. Especially at high FLOP counts, our laws predict overtraining becomes quite efficient in that it results in fairly small elevations in loss for a relatively large reduction in model size.
@@
Industry models often use fewer parameters and train on more tokens than prescribed in prior work. We find the impact of overtraining a smaller model on predicted loss to be small. Combining this with Figure˜16, where wider models are predicted to be optimal in terms of GPU hours, reinforces the message that FLOPs optimality is not the end of the story for training models. Trading some FLOPs optimality for time optimality necessarily means overtraining, but Figure˜17 suggests the difference is marginal. We believe this combined evidence makes significant progress towards explaining the differences between the prescriptions found in prior work and training choices observed in industry.
@@
J.2 Additional Training Complications
@@
Appendix K Ablations for Approach 1
K.1 Extended Paper Figures
In Figure˜19, we plot an extended version of the Approach 1 plot we present in Figure˜3.
@@
K.2 Alternative Learning Rates
In Figure˜20, we present the Approach 1 prescription when fitting on the learning rate ablation data.
@@
K.3 Cooldown
In Figure˜21, we present the Approach 1 prescription when fitting on the cooldown ablation data.
@@
K.4 Varying Delta in the Huber loss
In Figure˜22, where we plot the exponents found by optimizing the Huber loss versus the size of the grid search used for optimization. We see that a delta of $10^{-5}$ is unstable for smaller grid sizes and including more tokens in the fitting data generally increases stability of the exponents found during optimization.
@@
Appendix L FLOP counting matters

### v3

@@ Appendix I The Price of Stepping Off the Scaling Law
If You Value Your Time, Train Wide Models. We first show that in our training setup, training wider models is far more efficient than training deep models. In Figure˜17, we reflect on the consequences of suboptimal architectural choices, by considering how much of a given resource—FLOPs or GPU hours—would be “overspent” to reach any target loss value with the plotted architecture rather than the prescribed width and depth. We find that choosing to train “skinny” models (top left) wastes many FLOPs and GPU hours. The scale of overspend is quite different however, with the least efficient models only overspending about $50\%$ on FLOPs but wasting more than $200\%$ of the GPU hours spent by the best configuration. In other words, in the time taken to train a single (very) suboptimal model to the desired loss value, one could train three optimal-width-depth models. We note that while the time-optimal models tend to be the wider ones, this is probably due to our training scheme. Similar to other open-source efforts such as OLMo et al. [2024], we do not make any use of pipeline parallelism. In summary, for standard parallelism implementations, wider models are simply easier to scale, but as a result our observations regarding resource overspending may not generalize to other parallelism strategies.
Appendix J Scaling Laws Predict That Overtraining Is Efficient.
@@
Similarly to Gadre et al. [2024], we can shift optimal points to simulate overtraining. To do this, we fix a FLOP budget and trace out a path of model sizes and corresponding token counts to remain within that budget. For each model size and token count, we record the “overtraining factor,” which is the selected number of training tokens divided by the optimal number of tokens for that model shape. An overtraining factor of less than one corresponds to undertraining the model, and a factor greater than one represents overtraining. We show the results of this process in Figure˜18. We see that overtraining does increase predicted loss at a given FLOP count but that these curves are actually quite flat. We include the loss values of open source models on our own validation set to allow readers to contextualize the y-axis values. Especially at high FLOP counts, our laws predict overtraining becomes quite efficient in that it results in fairly small elevations in loss for a relatively large reduction in model size.
@@
Industry models often use fewer parameters and train on more tokens than prescribed in prior work. We find the impact of overtraining a smaller model on predicted loss to be small. Combining this with Figure˜17, where wider models are predicted to be optimal in terms of GPU hours, reinforces the message that FLOPs optimality is not the end of the story for training models. Trading some FLOPs optimality for time optimality necessarily means overtraining, but Figure˜18 suggests the difference is marginal. We believe this combined evidence makes significant progress towards explaining the differences between the prescriptions found in prior work and training choices observed in industry.
@@
K.2 Additional Training Complications
@@
Appendix L Ablations for Approach 1
L.1 Extended Paper Figures
In Figure˜20, we plot an extended version of the Approach 1 plot we present in Figure˜3.
@@
L.2 Alternative Learning Rates
In Figure˜21, we present the Approach 1 prescription when fitting on the learning rate ablation data.
@@
L.3 Cooldown
In Figure˜22, we present the Approach 1 prescription when fitting on the cooldown ablation data.
@@
L.4 Varying Delta in the Huber loss
In Figure˜23, where we plot the exponents found by optimizing the Huber loss versus the size of the grid search used for optimization. We see that a delta of $10^{-5}$ is unstable for smaller grid sizes and including more tokens in the fitting data generally increases stability of the exponents found during optimization.
@@
Appendix M FLOP counting matters